Find the general solution to the differential equation: ty'+2y =t3, and what are the homogenous solutions. Find the general solution to the differential equation: ty+2y = 3
Answer #1

tytzy=t3 y + 2 y = t² + S 2 / dt e Integrating factor - ent 희은 a e t solution of ( y.+?)- leit tc (y +2) 15 f С 3 t5 so + ct-ty + 2y = 0 ty =-zy 2 bete t 2 lut + luc luy lug-luc = lu ( 7 / 2 y = ceta Homogtious aduation